OpenStudy (anonymous):

solve by factoring: x^2=14x-33

OpenStudy (anonymous):

move everything to one side first

OpenStudy (anonymous):

14x-33-x^2?

OpenStudy (anonymous):

is equal to 0 so 14x-33-x^2=0 preferably we want x^2 to be positive, so multiply by -1 or move everythign to the otherside will give x^2-14x+33=0

OpenStudy (anonymous):

now, solving by factoring, list all the factor couples of 33

OpenStudy (anonymous):

11

OpenStudy (anonymous):

11x3

OpenStudy (anonymous):

factor pairs, multiple pairs, same thing ex. [1,33], [3,11], etc

OpenStudy (anonymous):

there should be two more left, list them

OpenStudy (anonymous):

1x33

OpenStudy (anonymous):

i already listed that one,

OpenStudy (anonymous):

thats it. 11x3 and 1x33

OpenStudy (anonymous):

nope -1 *-33 =?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

(x-11) (x-3)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well, almost, you forgot =0

OpenStudy (anonymous):

(x-11) (x-3)=0. thanks so much!

OpenStudy (anonymous):

not done yet

OpenStudy (anonymous):

oh?

OpenStudy (anonymous):

zero property of multiplication means that either (x-11) =0 or (x-3)=0 so basically just solve for x

OpenStudy (anonymous):

so x will have 2 answers

OpenStudy (anonymous):

11, 3?

OpenStudy (anonymous):

x= 11, x =3 its an equation, thus the answer would be an equation

OpenStudy (anonymous):

but yes x is equal to 11 and 3

OpenStudy (anonymous):

ok. i dont have to put it in an equation. i just had to put the two x values. so thanks for helping :)

OpenStudy (anonymous):

computer assignment?

OpenStudy (anonymous):

some teachers will be picky about how the answer is written

OpenStudy (anonymous):

yeah. and i know. i hate it but you helped a lot.

OpenStudy (anonymous):

no prob